Two campers dock a canoe. One camper steps onto the dock.  This camper has a mass of 80 kg and moves forward at 4.0 m/s. With what speed and direction do the canoe and the other camper move if their combined mass is 110 kg?


1
Expert's answer
2021-05-05T10:54:47-0400

According to the momentum conservation law, the momentum "p_1" of the first camper is equal to the momentum "p_2" of the boat with the second camper. By definition, the momentums are given as follows:


"p_1 = m_1v_1\\
p_2 = m_2v_2"

where "m_1 = 80kg, m_2 = 110kg" are the masses of the first camper, and the boat with the second camper respectively, "v_1 = 4m\/s,v_2" are the speeds of the first camper, and the boat with the second camper respectively.

Thus, obtain the equality:


"m_1v_1 = m_2v_2\\
v_2 = \dfrac{m_1v_1 }{m_2} = \dfrac{80kg\cdot 4m/s}{110kg} \approx 2.9m/s"

The direction of this velocity is opposite to the direction of "v_1" (thus, backward).


Answer. 2.9 m/s backward.